OpenStudy (anonymous):

Tell whether the expression is a polynomial. If it is, give its degree. -2\pi Seleccione una respuesta. a. polynomial, degree 0 b. polynomial, degree 1 c. not a polynomial d. polynomial, degree -2

OpenStudy (anonymous):

according to my old memory, polynomial is something related to an unknown like x, t, or y.... in this case -2/pi= it is actually a number..... so i dont think it is a polynomial thus, the ans is c

OpenStudy (anonymous):

The answer a is not right... what do you think about the answer answer....`?

OpenStudy (anonymous):

The d -2 is not right i am sure

OpenStudy (anonymous):

it should be b then coz x^0=1= constant..... it can be expressed as -2/pi*x^0 ....sr my bad

OpenStudy (anonymous):

Are you sure is grade 1?

OpenStudy (anonymous):

i mean a--degree 0

OpenStudy (anonymous):

sr. too tired....

OpenStudy (anonymous):

btw, good catch!!!!!!

OpenStudy (anonymous):

-2\pi ? is it -2/pi ?

OpenStudy (anonymous):

Iam sorry but -2 is wrong

OpenStudy (anonymous):

i think typing a polynomial is harder than finding if it is a polynomial :)